OpenStudy (anonymous):

Use the derivative rules to differentiate each of the following: 1. f(x)=1/x-1 and 2. f(x)=sqrt(x)

OpenStudy (anonymous):

still not clear what is meant by "rules" second one you should memorize \[f(x)=\sqrt{x},f'(x)=\frac{1}{2\sqrt{x}}\]

OpenStudy (anonymous):

(1.) Rewrite the function as an exponent so that you can use the power rule, but you will also have to use the chain rule. \[f(x)=1/(x+1) \]\[f(x)=(x-1)^{-1}\] Chain Rule: \[f \prime(x)= -(x-1)(1) = 1-x\] (2.) Write the square root function as x to the one half, again so that you can use the power rule. Remember that you must subtract 1 from the exponent (rename as 2/2) to get -1/2. \[ f(x)=\sqrt{x}\] \[f(x)=x^{\frac{ 1 }{ 2 }}\] \[f \prime(x)=\frac{ 1 }{ 2 } (x)^{\frac{ -1 }{ 2 }}\] \[=\frac{ 1 }{ 2 \sqrt{x}}\]

OpenStudy (anonymous):

ooops

OpenStudy (anonymous):

try the first one again

OpenStudy (anonymous):

Snap, I don't feel like writing it out again, but it is \[-(x+1)^{-2}\] which is equal to \[-\frac{ 1 }{( x+1)^{2} }\] Subtract 1 ! Sorry about that.